OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x - 6y = 12 -2x + 3y = 6

OpenStudy (anonymous):

do you know matrices?

OpenStudy (anonymous):

not to well my teacher put me in a class i have no idea about

OpenStudy (anonymous):

but have you ever experienced matrices? like heard of it?

OpenStudy (anonymous):

i heard of it never experienced it but someone on here tried to help but didnt work to good

OpenStudy (anonymous):

there is another method, which is just subsitution, its more basic but will still yield the same solution

OpenStudy (anonymous):

which would you prefer?

OpenStudy (anonymous):

the easiest one

OpenStudy (anonymous):

haha. alright lets not do matrices then. k lets re-arrange one equation, preferably the first equation as either x or y as the subject. can you do this?

OpenStudy (anonymous):

some what you might want to help

OpenStudy (anonymous):

Lets re-arrange the first as x as the subject alright. So we have 3x-6y=12 3x=6y+12 x=(6y+12)/3

OpenStudy (anonymous):

follow?

OpenStudy (anonymous):

multiply the second eq with 2 then add byt his u can cancel y and u will get the value of x substitute the obtained value of x in any one eq u will get the value of y

OpenStudy (anonymous):

x=2y+4

OpenStudy (anonymous):

yeah i follow

OpenStudy (anonymous):

alright. so now we have x=2y+4 and 3y=2x+6

OpenStudy (anonymous):

Simply substitute x=2y+4 into 3y=2x+6

OpenStudy (anonymous):

So we get: 3y=2(2y+4)+6

OpenStudy (anonymous):

get me?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

So now we just expand the brackets. We solve simultaneously because we now have an equation with one vairable (y). So: 3y=4y+8+6 3y-4y=14 -y=14 y=-14

OpenStudy (anonymous):

So we know y=-14 and that x=2y+4 Now we can find x by substituting y=-14 into that equation x=2(-14)+4 x=-28+4 x=-24

OpenStudy (anonymous):

therefore x=-24 y=-14 is your solution. That is, the two lines intersect at the point (-24,-14)

OpenStudy (anonymous):

thank you all for helping me because i get freaking lost when it come to math so i get rlly lost when it come to algebra

OpenStudy (anonymous):

Just be logical, follow the rules and PRACTICE!